rule in the following function
0 2 4 6
3 4 5 6

Answer by stanbon(75887)

rule in the following function
0 2 4 6
3 4 5 6
----------------
As x increases by 2, y increases by 1 so the slope is 1/2
---
Since 3 is matched up with 0 the intercept is 3
------------
Equation:
y = (1/2)x + 3
====================
Using the points you would get:
slope = (4-3)/(2-0) = 1/2
---
Since y = mx + b, 3 = (1/2)*0 + b; b = 3
Equation:
y = (1/2)x + 3
=========================
Cheers,
Stan H.